# Recovering a phase transition signal in simulated LISA data with a modulated galactic foreground

### v1

@@ 1 Introduction
The creation of catalogues of merging compact binaries by the LIGO, VIRGO, and KAGRA collaborations (LVK) [1] has solidified gravitational waves as a unique probe of the cosmos. In addition to these compact binaries, there are many cosmological and astrophysical sources that can produce a stochastic gravitational wave background (SGWB, see e.g. [2, 3]). Recent pulsar timing array data from NANOGrav [4], the Chinese Pulsar Timing Array [5], the European Pulsar Timing Array and Indian Pulsar Timing Array [6] and the Parkes Pulsar Timing Array [7] (see also Ref. [8] for a reanalysis of this data) show evidence for a background in the nHz range, although as yet there are only upper limits at LVK’s peak sensitivity of around 100 Hz [9]. Searching for such a background in the mHz frequency range is one of the science targets of the upcoming Laser Interferometer Space Antenna (LISA) mission [10], a space-based GW interferometer set to launch next decade with sensitivity to gravitational waves with frequencies between $10^{-4}$ Hz and $10^{-1}$ Hz [11].
@@
This paper is organised as follows. We begin by discussing the production and processing of simulated LISA data in Sections 2 and 3, respectively. In Section 4 we review the set-up of our MCMC analyses, before discussing our results in Section 5 and concluding in Section 6. Finally, in Appendix A we present the necessary steps to reproduce our simulated data; while in Appendix B we discuss the choices we made for the Fourier expansion used for the annual modulation; and in Appendix C we provide tables with all of the numerical values for our Figs. 3 and 4.
@@ 2 Simulated LISA data
In the following, we introduce our GW sources and the signals we expect from them. We also elaborate on the detector response and noise functions, writing out the analytical expressions to be used in our data analysis. Lastly, we outline the TDI calculations leading to laser noise cancellation. For a more detailed description of our simulation for data reproduction purposes, refer to Appendix A.
@@
Rather than simulating each of these binaries individually, which would be computationally expensive, we instead create a skymap to which we insert a function describing the confusion noise coming from the binary population. We justify this choice in Appendix A. There are various analytical functions obtained by fits to population models [57, 42, 58], where the confusion noise foreground has been shown to peak at around 2 mHz, with estimations for the peak amplitude ranging from $10^{-46}$ to $10^{-44}\;\mathrm{Hz}^{-1}$. In order to reduce the number of parameters while still maintaining a close fit to the binary data in the frequency range of interest, we choose the model from Ref. [58]:
@@
where $A_{\mathrm{gb}}=1.4\times 10^{-45}\;\mathrm{Hz}^{-1}$ is the peak amplitude, $f_{\mathrm{gb}}=1.29$ mHz is the peak frequency, $\alpha_{\mathrm{gb}}=-7/3$ is the spectral slope, and $\gamma_{\mathrm{gb}}=900$ is a fit parameter that depends on the properties of the GB population. Here we assume that the spectral shape is the same for all binary populations across the sky.
@@
LISA’s response to gravitational waves is anisotropic: waves incident from directions normal to the plane of the spacecraft constellation induce a larger response than those arriving in the plane. The plane of the constellation is inclined with respect to the plane of LISA’s orbit around the sun, and hence the response to a source with fixed sky position is modulated with period of one year. Galactic binaries are concentrated near the galactic plane, and therefore the amplitude of the GB confusion noise will naturally be modulated, and can be represented in Fourier modes with period $t_{\text{yr}}$, the sidereal year. The broad beam of the response function sweeps across the galaxy twice per year in different sky locations, and hence we expect the principal modulation to be in the first and second harmonics. Our simulated GB signal (see Figs. 1 and 2) clearly shows this effect. We therefore construct our model for the modulated GB confusion noise as
@@
where $\omega_{n}=2\pi n/t_{\text{yr}}$. We investigate the effect of including higher harmonics in Appendix B. We find that their inclusion does not significantly improve the ability to resolve a PT signal.
@@
Time-delay interferometry
@@
At low frequencies, the $X_{2}$ transfer function for gravitational waves behaves as $(f/f_{*})^{6}$. As for the instrument noise, again assuming constant and equal arm length, our two noise components from Eqs. 2.7 and 2.8 enter the full noise power spectral density for TDI $X_{2}$ as
@@ 3 Data processing
with a 50% overlap [94, 95]. This compensates for the loss of time-domain data resulting from the attenuation of the signal at the start and end of each segment, as every data point is effectively counted twice. The Welch estimate of the PSDs is carried out using the welch function in the SciPy signal processing toolbox [96].
@@
Figure 2: Total power spectral density of the full dataset (black) and of each chunk (colour gradient). The grey hatched regions show the frequency limits for the MCMCs on both the chunked and unchunked data.
@@
4 Model comparison and statistical tests
We simulate data from LISA for a duration of $T_{\text{obs}}$ with a sample interval $T_{\text{samp}}$, yielding a time series of $L=T_{\text{obs}}/T_{\text{samp}}$ data points $X_{2}(t)$. For now, we consider the simplest case, meaning data without chunking, without overlaps in the segments, and without applying a window function to the segments. This corresponds to the method of averaged periodograms, also known as Bartlett’s method. We therefore start with $N_{\mathrm{w}}$ non-overlapping segments containing $L_{\text{w}}=L/N_{\mathrm{w}}$ data points, and assume that $L_{\text{w}}$ is an integer. The Fourier transform of the $i$th segment $\tilde{X}^{i}_{2}(f_{n})$ contains frequencies $f_{n}=n/L_{\text{w}}T_{\text{samp}}$ with $0\leq n<L_{\text{w}}$. Frequencies with $L_{\text{w}}/2<n$ are complex conjugates of those with $L_{\text{w}}-n$, hence there are $N_{\mathrm{f}}=L_{\text{w}}/2$ independent frequencies. The one-side power spectral density $P^{i}(f_{n})$ of the $i$th segment is
@@ 4 Model comparison and statistical tests
Welch’s method improves on the method of averaged periodograms by introducing a window function to reduce spectral leakage, and overlaps to compensate for the loss of information. As discussed in Section 3, we use a Hann window with a $50\%$ overlap. With windows of length $L_{\text{w}}$ and a 50% overlap, the number of windows is $N_{\text{w}}=2L/L_{\text{w}}-1$. The overlaps introduce correlations between $P^{i}(f_{n})$ from different segments; nevertheless the probability distribution for the mean $\overline{P}(f_{n})$ is well approximated by a chi-squared distribution, with an adjusted degrees of freedom parameter. To obtain the degrees of freedom in this case, we follow the calculation done in Refs. [92, 91] [^2] We note that in the last equation on page 29 of Ref. [91] there is a missing factor 2 in the denominator, which does appear in the similar equation 292b of Ref. [92], which gives
@@
Additionally, as discussed in Section 3, we only consider a subset $F$ of the full set of frequencies, so that the total number of frequencies is less than $L_{\text{w}}$. Therefore, our final likelihood function will be given by Eq. (4.3) with the product taken over a subset of frequencies $f_{n}\in F$, with ${N}_{\mathrm{w}}=2L/L_{\text{w}}-1$, and $\nu$ as given in Eq. (4.4).
@@
where $\{S_{k}\}$ is the set of model PSDs, which includes annual modulation.
Once we have the simulated and processed data, we use the parameter extraction code cobaya [98] to recover both the injected noise contributions and the injected PT signal via MCMC analyses.
$\text{DIC}=D(\overline{\theta})+2p_{D}\,.$ (4.6)
@@ 0 No No No
All of our analyses will feature the following base parameters:
  • Three parameters describing the white dwarf confusion noise: $A_{\mathrm{gb}}$, $f_{\mathrm{gb}}$, and $\gamma_{\mathrm{gb}}$. We fix $\alpha_{\mathrm{gb}}=-7/3$ in all our analyses.
These parameters will be kept fixed across all of our simulated datasets [^3] While the settings and input parameters for the different noise sources are the same, the generated data will be different in each run, as the noise itself is generated from a random seed; however, these noise contributions will share the same statistical properties.. All of our datasets will also include an injected PT signal, given by a combination of $\Omega_{\mathrm{pt}}$ and $f_{\mathrm{pt}}$. Specifically, we will consider all combinations of
[…]
When running our MCMCs, we will always include the base parameters, and we will have the following parameters that will be switched on or off in our different analyses:
  • Four parameters describing the annual modulation of the galaxies: two amplitudes $A_{1}$ and $A_{2}$, and two phases $\phi_{1}$ and $\phi_{2}$, as seen in equation 2.2. See Appendix B for more information about this choice.
@@ Acknowledgements
We thank Jean-Baptiste Bayle for insightful discussions in the early stages of this work. MH (ORCID ID $0$-0002-9307-437X) was supported by Academy of Finland grant no 333609. DCH (ORCID ID $0$-0003-2811-0917) was supported by Academy of Finland grant nos. 328958, 349865 and 353131. TM (ORCID ID 0009-0005-6855-184X) was supported by Academy of Finland grant no. 349865 and the Magnus Ehrnrooth Foundation. DJW (ORCID ID $0$-0001-6986-0517) was supported by Academy of Finland grant nos. 324882, 328958, 349865 and 353131. We acknowledge CSC – IT Center for Science, Finland, for computational resources.
@@ Appendix A Simulating LISA data
In order to simulate the galactic binaries, we create a skymap with HEALPix, into which we insert the confusion noise coming from the binary population. The angular noise power spectrum for LISA is expected to increase substantially for multipoles $\ell>6$, as seen in Fig. 13 of Ref. [61] for one year of data and Fig. 4 of Ref [105] for four years of LISA data. We can convert this multipole to a corresponding solid angle via $\Omega=2(180\deg)^{2}/(\pi\ell)$. Following Table 1 of Ref. [106], for this solid angle we need to take a HEALPix skymap with at least $N_{\mathrm{pix}}=48$ pixels. We therefore create a HEALPix skymap with 48 pixels, into which we insert the coordinates of the GBs contained in the catalogue, such that the skymap pixel intensity is defined by the number of GBs in the pixel normalised by the total number of GBs. Because LISA GW Response requires the square root of the intensity skymap, we further take the square root of our skymap. The resulting skymap is presented in the right panel of Fig. 7. A more precise distribution of the binaries entering our skymap is given in the left panel of the same figure.
@@ Appendix B Choosing the number of parameters for the annual modulation
To evaluate how many parameters we need to describe the annual modulation of the GB signal, we perform three tests. First, we investigate the amplitude of the real and imaginary Fourier coefficients, averaged over the 48 chunks. A similar analysis can be found in Fig. 3 in Ref [55]. We show the amplitude average for the first 10 coefficients in Fig. 8, where we can see that only the first two terms have a noticeable deviation from zero.
Figure 8: Amplitude of the first real and imaginary Fourier coefficients, averaged over the 48 chunks.
To investigate the difference between including two or three Fourier terms, we perform three MCMCs on data which only contains the GB signal, without any instrument noise or PT signal. In the first MCMC, we do not include any annual modulation parameters in the analysis (model Mc), in the second MCMC we include two terms of the Fourier expansion (M2c), and in the third MCMC we include three terms of the Fourier expansion (M3c). By comparing the $\Delta$DICs between these models, we can evaluate if the goodness-of-fit improves when including the various Fourier terms. Between models Mc and M2c, we find $\Delta\mathrm{DIC}\approx 10^{6}$, and between models M2c and M3c we find $\Delta\mathrm{DIC}\approx 10^{2}$. This shows that the inclusion of the first two Fourier terms provides a substantial improvement in the goodness-of-fit, while including the third term only moderately improves the fit further.

### v2

@@ 1 Introduction
The creation of catalogues of merging compact binaries by the LIGO, VIRGO, and KAGRA collaborations (LVK) [1] has solidified gravitational waves as a unique probe of the cosmos. In addition to these compact binaries, there are many cosmological and astrophysical sources that can produce a stochastic gravitational wave background (SGWB, see e.g. [2, 3]). Recent pulsar timing array data from NANOGrav [4], the Chinese Pulsar Timing Array [5], the European Pulsar Timing Array and Indian Pulsar Timing Array [6] and the Parkes Pulsar Timing Array [7] (see also Ref. [8] for a reanalysis of this data) show evidence for a background in the nHz range, although as yet there are only upper limits at LVK’s peak sensitivity of around 100 Hz [9]. Searching for such a background in the mHz frequency range is one of the science targets of the upcoming Laser Interferometer Space Antenna (LISA) mission [10], a space-based GW interferometer set to launch next decade with sensitivity to gravitational waves with frequencies between $10^{-4}$ Hz and $10^{-1}$ Hz [11, 12].
@@
This paper is organised as follows. We begin by discussing the production and processing of simulated LISA data in Sections 2 and 3, respectively. In Section 4 we derive our likelihood function and review the set-up of our MCMC analyses, before discussing our results in Section 5 and concluding in Section 6. Finally, in Appendix A we present the necessary steps to reproduce our simulated data; while in Appendix B we discuss the choices we made for the Fourier expansion used for the annual modulation; and in Appendix C we provide tables with all of the numerical values for our Figs. 3 and 4.
@@ 2 Simulated LISA data
In the following, we introduce our GW sources and the signals we expect from them. We also elaborate on the detector response and noise functions, writing out the analytical expressions to be used in our data analysis. Lastly, we outline the TDI calculations leading to laser noise cancellation. For a more detailed description of our simulation for data reproduction purposes, refer to Appendix A. Additionally, we have made our full data simulation and analysis pipeline publicly available.111zenodo.org/doi/10.5281/zenodo.12781278
@@
Rather than simulating each of these binaries individually, which would be computationally expensive, we instead create a skymap representing the spatial distribution of binaries in LDC2a. We then insert a function describing the confusion noise coming from the binary population into the pixels of the skymap. We justify and elaborate on this choice in Appendix A. There are various analytical functions obtained by fits to population models [58, 43, 59], where the confusion noise foreground has been shown to peak at around 2 mHz, with estimations for the peak amplitude ranging from $10^{-46}$ to $10^{-44}\;\mathrm{Hz}^{-1}$. These confusion noise models are for the unresolved GB, as GBs with high signal-to-noise ratio are expected to be individually resolvable by LISA, and their inclusion would significantly enhance the overall signal [81, 82, 83]. In order to reduce the number of parameters while still maintaining a close fit to the unresolved GB data in the frequency range of interest, we choose the model from Ref. [59]:
@@
where $A_{\mathrm{gb}}=1.4\times 10^{-45}\;\mathrm{Hz}^{-1}$ is the peak amplitude, $f_{\mathrm{gb}}=1.29$ mHz is the peak frequency, $\alpha_{\mathrm{gb}}$ is the spectral slope, and $\gamma_{\mathrm{gb}}=900$ is a fit parameter that depends on the properties of the GB population. Here we assume that the spectral shape is the same for all binary populations across the sky. As in the aforementioned references, we fix $\alpha_{\mathrm{gb}}=-7/3$ as per the theoretical prediction for gravitational wave emission from a binary population [58].
@@
LISA’s response to gravitational waves is anisotropic: waves incident from directions normal to the plane of the spacecraft constellation induce a larger response than those arriving in the plane. The plane of the constellation is inclined with respect to the plane of LISA’s orbit around the sun, and hence the response to a source with fixed sky position is modulated with period of one year. Galactic binaries are concentrated near the galactic plane, and therefore the amplitude of the GB confusion noise will naturally be modulated, and can be represented in Fourier modes with period $t_{\text{yr}}$, the sidereal year:
$S_{\text{gb}}(f,t)=S_{\mathrm{gb}}(f)\left[1+\sum_{n=0}^{n}A_{n}\cos(\omega_{n}t+\phi_{n})\right]\,,$ (2.2)
where $\omega_{n}=2\pi n/t_{\text{yr}}$ and $A_{n}$ are the Fourier coefficients (for a full derivation of these, see e.g. [84]). The broad beam of the response function sweeps across the galaxy twice per year in different sky locations, and hence we expect the principal modulation to be in the first and second harmonics, as shown in e.g. [84, 85]. Our simulated GB signal (see Fig. 1) clearly shows this effect. We therefore construct our model for the modulated GB confusion noise as
@@
We investigate the effect of including higher harmonics in Appendix B. We find that their inclusion does not significantly contribute to the GB modulation, and does not improve the ability to resolve a PT signal.
@@
Time-delay interferometry:
@@
At low frequencies, the $X_{2}$ transfer function for gravitational waves behaves as $(f/f_{*})^{6}$. As for the instrument noise, again assuming constant and equal arm length, our two noise components from Eqs. 2.8 and 2.9 enter the full noise power spectral density for TDI $X_{2}$ as
@@ 3 Data processing
with a 50% overlap [100, 101]. This compensates for the loss of time-domain data resulting from the attenuation of the signal at the start and end of each segment, as every data point is effectively counted twice. The Welch estimate of the power spectral densities is carried out using the welch function in the SciPy signal processing toolbox [102].
@@
Figure 2: Total power spectral density of the full dataset (black) and of each chunk (colour gradient). The grey hatched regions show the frequency limits for the MCMCs on both the chunked and unchunked data. The bottom panel shows a zoomed in version of the region where the annual modulation has the biggest impact.
@@
4 Likelihood and model comparison
4.1 Likelihood for averaged spectra
General case:
Consider a time series of duration $T_{\text{obs}}$ with a sample interval $T_{\text{samp}}$, yielding a time series of $L=T_{\text{obs}}/T_{\text{samp}}$ data points $X_{2}(t)$. For now, we consider the simplest case, meaning data without chunking, without overlaps in the segments, and without applying a window function to the segments. This corresponds to the method of averaged periodograms, also known as Bartlett’s method. We therefore start with $N_{\mathrm{w}}$ non-overlapping segments containing $L_{\text{w}}=L/N_{\mathrm{w}}$ data points, and assume that $L_{\text{w}}$ is an integer. The Fourier transform of the $i$th segment $\tilde{X}^{i}_{2}(f_{n})$ contains frequencies $f_{n}=n/L_{\text{w}}T_{\text{samp}}$ with $0\leq n<L_{\text{w}}$. Frequencies with $L_{\text{w}}/2<n$ are complex conjugates of those with $L_{\text{w}}-n$, hence there are $N_{\mathrm{f}}=L_{\text{w}}/2$ independent frequencies. The one-side power spectral density $P^{i}(f_{n})$ of the $i$th segment is
@@ 4.1 Likelihood for averaged spectra
Specialising to overlapping windowed data:
Welch’s method improves on the method of averaged periodograms by introducing a window function to reduce spectral leakage, and overlaps to compensate for the loss of information. When using Welch’s method and a Hann window with a $50\%$ overlap and with windows of length $L_{\text{w}}$, the number of windows is $N_{\text{w}}=2L/L_{\text{w}}-1$. The overlaps introduce correlations between $P^{i}(f_{n})$ from different segments; nevertheless the probability distribution for the mean $\overline{P}(f_{n})$ is well approximated by a chi-squared distribution, with an adjusted degrees of freedom parameter. To obtain the degrees of freedom for this case, we follow the calculation done in Refs. [98, 97] [^3] We note that in the last equation on page 29 of Ref. [97] there is a missing factor 2 in the denominator, which does appear in the similar equation 292b of Ref. [98]., which gives
@@
4.2 MCMC methodology for modulated LISA data
We now apply the above general likelihood to the case of our simulated and processed LISA data. As discussed in Section 3, we consider a subset $F$ of the full set of frequencies, so that the total number of frequencies is less than $L_{\text{w}}$. Therefore, our final likelihood function will be given by Eq. (4.3) with the product taken over a subset of frequencies $f_{n}\in F$, with ${N}_{\mathrm{w}}=2L/L_{\text{w}}-1$, and $\nu$ as given in Eq. (4.4).
@@ 4.2 MCMC methodology for modulated LISA data
where $\{S_{k}\}$ is the set of model power spectral densities, which includes annual modulation.
We implement these likelihood functions in the parameter extraction code cobaya [104], which we use to recover both the injected noise contributions and the injected PT signal via MCMC analyses. Specifically, we use the mcmc sampler in cobaya, which is based on the Metropolis Hastings algorithm as presented in [105]. As we do not encounter any multi-modal distributions in our results, this sampler is sufficient for our analyses.
All of our analyses will feature the following base parameters:
  • Three parameters describing the white dwarf confusion noise: $A_{\mathrm{gb}}$, $f_{\mathrm{gb}}$, and $\gamma_{\mathrm{gb}}$. We fix $\alpha_{\mathrm{gb}}=-7/3$ in all our analyses as per the discussion in Section 2.
These parameters will be kept constant in all of our simulated datasets. [^4] While the settings and input parameters for the different noise sources are the same, the generated data will be different in each run, as the noise itself is generated from a random seed; however, these noise contributions will share the same statistical properties. All of our datasets will also include an injected PT signal, given by a combination of $\Omega_{\mathrm{pt}}$ and $f_{\mathrm{pt}}$. Specifically, we will consider all combinations of
[…]
When running our MCMCs, we will always vary the base parameters, and we will have the following parameters that will be either varied in the MCMC analyses or set to 00 in our different analyses:
  • Four parameters describing the annual modulation of the galaxies: two amplitudes $A_{1}$ and $A_{2}$, and two phases $\phi_{1}$ and $\phi_{2}$, as seen in Eq. 2.3. See Appendix B for more information about this choice.
For all parameters, we choose the starting values for the MCMC analyses to be the values injected in the simulated data, and we chose the prior range so that the distributions are approximately symmetric around the starting values. The exact prior ranges for the parameters are listed in Table 1.
Prior Starting value Minimum Maximum
$\log_{10}S_{\mathrm{a}}$ logarithmic -29.05 -23.24 -34.85
$\log_{10}S_{\mathrm{p}}$ logarithmic -21.65 -17.32 -25.98
$\log_{10}A_{\mathrm{gb}}$ logarithmic -44.85 -48 -42
$\log_{10}f_{\mathrm{gb}}$ logarithmic -2.89 -3.30 -2.52
$\gamma_{\mathrm{gb}}$ linear 900 500 1200
$\log_{10}\Omega_{\mathrm{pt}}$ logarithmic model specific starting value -1 starting value +1
$\log_{10}f_{\mathrm{pt}}$ logarithmic model specific starting value -1 starting value +1
$A_{1}$ linear 0.1 0 1
$A_{2}$ linear 0.5 0 1
[…]
Table 1: Prior type, starting value, and prior range for all of the parameters that can be varied in our different MCMC analyses.
@@ Acknowledgements
We thank Jean-Baptiste Bayle for insightful discussions and feedback. MH (ORCID ID $0$-0002-9307-437X) was supported by Academy of Finland grant no 333609. DCH (ORCID ID $0$-0003-2811-0917) was supported by Academy of Finland grant nos. 328958, 349865 and 353131. TM (ORCID ID 0009-0005-6855-184X) was supported by Academy of Finland grant no. 349865 and the Magnus Ehrnrooth Foundation. DJW (ORCID ID $0$-0001-6986-0517) was supported by Academy of Finland grant nos. 324882, 328958, 349865 and 353131. We acknowledge CSC – IT Center for Science, Finland, for computational resources.
@@ Appendix A Simulating LISA data
In order to simulate the galactic binaries, we start by creating a HEALPix skymap. The angular noise power spectrum for LISA is expected to increase substantially for multipoles $\ell>6$, as seen in Fig. 13 of Ref. [62] for one year of data and Fig. 4 of Ref [112] for four years of LISA data. We can convert this multipole to a corresponding solid angle via $\Omega=2(180\deg)^{2}/(\pi\ell)$. Following Table 1 of Ref. [113], for the solid angle corresponding to $\ell>6$ we would need to take a HEALPix skymap with $N_{\mathrm{pix}}>12$ pixels. We therefore create a skymap with 48 pixels, into which we insert the coordinates of the GBs contained in the catalogue, such that the skymap pixel intensity is defined by the number of GBs in the pixel normalised by the total number of GBs. Because LISA GW Response requires the square root of the intensity skymap, we further take the square root of our skymap. The resulting skymap is presented in the right panel of Fig. 7. A more precise distribution of the binaries entering our skymap is given in the left panel of the same figure.
@@ Appendix B Choosing the number of parameters for the annual modulation
Figure 8: The first 10 real and imaginary Fourier coefficients $A_{n}\cos(\phi_{n})$ and $A_{n}\sin(\phi_{n})$ representing the modulation of the GB foreground, as given in Eq. 2.2.
First we wish to evaluate how many parameters we need to describe the annual modulation of the GB signal. To do this, we produce a dataset which only contains the GB component, i.e. is free of instrument noise and PT signals, and which can therefore be described by Eq. 2.1. As described in Section 3, to account for the annual modulation we divide our data into $N_{\mathrm{ch}}=48$ time chunks, and we approximate the amplitude of the power spectral density as being constant in each chunk. We then apply the WOSA method to obtain the power spectral density in each chunk, as described in Eq. 4.2.
Finally, we perform a discrete Fourier transform on the average power spectral density of each chunk (as shown in Eq. 2.2) to assess the contribution of each Fourier mode to the total GB signal. We show the amplitude of the real and imaginary Fourier coefficients in Fig. 8, where we can see that only the first two terms have a noticeable deviation from zero. Similar analyses can be found in [114, 56].
To investigate the statistical difference between including two or three Fourier terms, we perform three MCMCs on the same GB-only data. In the first MCMC, we do not include any annual modulation parameters in the analysis (model Mc), in the second MCMC we include two terms of the Fourier expansion (M2c), and in the third MCMC we include three terms of the Fourier expansion (M3c). By comparing the $\Delta$DICs between these models, we can evaluate if the goodness-of-fit improves when including the various Fourier terms. Between models Mc and M2c, we find $\Delta\mathrm{DIC}\approx 10^{6}$, and between models M2c and M3c we find $\Delta\mathrm{DIC}\approx 10^{2}$. This shows that the inclusion of the first two Fourier terms provides a substantial improvement in the goodness-of-fit, while including the third term only moderately improves the fit further.